OpenStudy (me123):

solve the compound inequality 5>-3x+4 or 10≤-5x+4

OpenStudy (anonymous):

Does the following look about right? \[x\leq -\frac{6} {5}\|x>-\frac{1}{3} \]

OpenStudy (anonymous):

Does the following look about right? \[x\leq -\frac{6} {5}\|x>-\frac{1}{3} \]

OpenStudy (me123):

yes

OpenStudy (me123):

is it infinity -1/3 or -1/3 infinity

OpenStudy (anonymous):

Well the result is from executing Reduce[5>-3x+4 or 10≤-5x+4,x] in Mathematica. Have not solved inequalities manually for decades. Not sure where infinity figures into this problem. Is this your question or did someone else put you up to it?

OpenStudy (me123):

this is my question from algebra

OpenStudy (me123):

it is multiple choice and one of the answers looks like this (00-5/6) upside down u (-1/3 -00)

OpenStudy (me123):

the 00 are infinity signs

OpenStudy (anonymous):

Looking back at the answer I would guess that -1/3 infinity is the answer because x has to be greater than -.033333 all the way to infinity if desired.

OpenStudy (me123):

Thank you

OpenStudy (anonymous):

Thank you.